# Theory of excitonic order in kagome metal ScV6Sn6

### v2

@@ Acknowledgements
In this section we discuss the mean-field analysis for excitonic order from short-ranged interactions. The interactions which generate excitonic order are those which couple the opposite concavity vHS, which are purely composed of the $d_{z^{2}}$ and $d_{xy}/d_{x^{2}-y^{2}}$ orbitals respectively. The interactions which drive excitonic order are therefore strictly those which couple the two distinct orbital types, which we index by $\alpha$ as in the main text. We write down an extended Hubbard model featuring only couplings between different orbitals, and between both electrons onsite and on neighbouring sites i.e. adjacent sublattices indexed by $\sigma$. We will use slightly different notation to the main text to better compare with prior work Scammell2023bib ; the $U^{\prime}$ and $V^{\prime}$ of the main text correspond to $U_{h_{4}}$ and $U_{l_{2}}$ below. The result is an interacting Hamiltonian,
@@
Figure S3: Phase diagram as a function of Hubbard parameters. Plot of the state with largest eigenvalue of the gap equation as a function of the Hubbard parameters; all units of energy are in eV. In the notation of the main text, $U_{l_{2}}=V^{\prime}$ and $U_{h_{4}}=U^{\prime}$ (a) $U_{h_{1}}$ versus $U_{l_{4}}$ at fixed $U_{h_{4}}=1$, with $U_{l_{2}}=U_{l_{4}}$ and $U_{j_{1}}=U_{j_{2}}=U_{j_{4}}=0$. (b) Setting $U_{h_{4}}=1$, and setting all $U_{l_{i}}=U_{h_{1}}$ versus $U_{j_{i}}$. (c) Taking values $U_{l_{2}}=U_{l_{4}}=0.5$, $U_{h_{4}}=1$, and comparing $U_{j_{i}}$ against $U_{h_{1}}$. (d) Taking $U_{h_{4}}=1,U_{h_{1}}=0.75$, setting all $U_{j_{i}}$ equal and $U_{l_{i}}$ equal and plotting $U_{l_{i}}$ against $U_{j_{i}}$.

### v3

@@ Acknowledgments
S1.2 Electronic structure of LuNb6Sn6
In this subection, we present first principles results for LuNb6Sn6, showing that the physics of ScV6Sn6 applies in this system as well. In Fig. S2, we illustrate the $k_{z}$ dispersion within the CDW phase; one observes the gapping out of the dispersion along the $\Gamma$-A line and depletion of $k_{z}$ spectral weight due to the out-of-plane distortion of the Lu and Sn atoms. As in ScV6Sn6, this justifies the use of a two-dimensional kagome model within the CDW phase. Fitting the bands to a tight-binding model of the form Eq. 1, we find the parameters $t_{x}=0.76$, $t_{z}=-0.30$, $\varepsilon_{x}=1.50$ and $\varepsilon_{z}=-0.77$ (all units in eV); the comparison between the two-orbital model and the DFT bandstructure is shown in Fig. S3, illustrating good agreement. As in the case of ScV6Sn6, we find that the two vHS near the Fermi level are $p$-type in nature, so that our arguments regarding the propensity to excitonic order carry through in this case as well.
Figure S2: Influence of CDW order on the out-of-plane dispersion in LuNb6Sn6. (a) Electronic structure of LuNb6Sn6 in the CDW phase, along the $\Gamma$-A path. (b) Electronic structure of LuNb6Sn6 in the CDW phase, along the A-H-L-A path. Note the opening of the Nb band around the Fermi level, similarly to the V band of ScV6Sn6 in the CDW configuration.
Figure S3: Electronic structure of LuNb6Sn6. (a) Unfolded band structure for LuNb6Sn6 in the CDW phase. (b) Comparison between Nb $d_{xy}$ + $d_{x^{2}-y^{2}}$ orbitally-projected DFT bands, in the pristine phase, and corresponding TB model. (c) Comparison between Nb $d_{z^{2}}$ orbitally-projected DFT bands, in the pristine phase, and corresponding TB model. (d) Comparison between Nb $d_{z^{2}}$ + $d_{xy}$ + $d_{x^{2}-y^{2}}$ orbitally-projected DFT bands, in the pristine phase, and the complete 12 bands TB model.
S1.3 Results for non-CDW 166 materials (Ho,Tb)Nb6Sn6
We contrast ScV6Sn6 and LuNb6Sn6 with other 166 kagome metals (Ho,Tb)Nb6Sn6, for which we find no CDW instability. Consequently, these systems have significant $k_{z}$ dispersion which is expected to suppress excitonic order, as $k_{z}^{2}$ enters the mean field excitonic Hamiltonian with the same form as a momentum-dependent Zeeman field in the Bogoliubov-de Gennes Hamiltonian for superconducting pairing. To describe the bandstructure of these materials we find it necessary to introduce $c$-axis hoppings $t^{(c)}_{i}$ with $i=x,z$; fitting tight-binding parameters for $A$Nb6Sn6, we find $\{\varepsilon_{x},\varepsilon_{z},t_{x},t_{z},t^{(c)}_{x},t^{(c)}_{z}\}=\{1.68,-0.61,0.76,-0.30,0.75,0.44\}$ for $A$=Ho and $\{\varepsilon_{x},\varepsilon_{z},t_{x},t_{z},t^{(c)}_{x},t^{(c)}_{z}\}=\{1.68,-0.61,0.76,-0.30,0.75,0.44\}$ for $A$=Tb. The significant $c$-axis dispersion in these systems distinguishes their physics from the CDW phase of Sc166 and Lu166.
Figure S4: Electronic structure of (Lu,Ho,Tb)Nb6Sn6. Comparison between Nb $d_{z^{2}}$ orbitally-projected DFT bands in the pristine phase, and corresponding TB model, for LuNb6Sn6 (a), HoNb6Sn6 (b) and TbNb6Sn6.
@@
In this section we discuss the mean-field analysis for excitonic order from short-ranged interactions. The interactions which generate excitonic order are those which couple the opposite concavity vHS, which are purely composed of the $d_{z^{2}}$ and $d_{xy}/d_{x^{2}-y^{2}}$ orbitals respectively  [^9] The vHS due to the $d_{xz}$/$d_{yz}$ lie a distance $\sim$ 0.25 eV away from the Fermi level, and and so do not contribute significantly to the Fermi pockets near $M$.. The interactions which drive excitonic order are therefore strictly those which couple the two distinct orbital types, which we index by $\alpha$ as in the main text. We write down an extended Hubbard model featuring only couplings between different orbitals, and between both electrons onsite and on neighbouring sites i.e. adjacent sublattices indexed by $\sigma$. We will use slightly different notation to the main text to better compare with prior work Scammell2023bib ; the $U^{\prime}$ and $V^{\prime}$ of the main text correspond to $U_{h_{4}}$ and $U_{l_{2}}$ below. The result is an interacting Hamiltonian,
@@
Figure S6: Phase diagram as a function of Hubbard parameters. Plot of the state with largest eigenvalue of the gap equation as a function of the Hubbard parameters; all units of energy are in eV. In the notation of the main text, $U_{l_{2}}=V^{\prime}$ and $U_{h_{4}}=U^{\prime}$. (a) $U_{h_{1}}$ versus $U_{l_{4}}$ at fixed $U_{h_{4}}=1$, with $U_{l_{2}}=U_{l_{4}}$ and $U_{j_{1}}=U_{j_{2}}=U_{j_{4}}=0.45$. (b) $U_{j_{i}}=U_{h_{1}}$ versus $U_{l_{i}}$, for all $i$, at fixed $U_{h_{4}}=1$. (c) $U_{j_{i}}$ versus $U_{h_{1}}$, for all $i$, at fixed $U_{l_{2}}=U_{l_{4}}=0.5,U_{h_{4}}=1$. (d) $U_{l_{i}}$ versus $U_{j_{i}}$, for all $i$, at fixed $U_{h_{4}}=1,U_{h_{1}}=0.75$.